A scuba diving resort hotel offers divers 2 plans. Plan A gives 3 nights of lodging and 4 dives. Plan B gives 5 nights of lodging and 8 dives. A price list comes out in which Plan A costs $440 and Plan B costs $780. What is the cost of a single dive?
Answer by TimothyLamb(4379)

x = dollar cost per dive
y = dollar cost per night
---
4x + 3y = 440
8x + 5y = 780
---
put the system of linear equations into standard form
---
4x + 3y = 440
8x + 5y = 780
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 35
y = 100
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---